OpenStudy (klepez):

can you walk me through 2(x+2)^2 +3

OpenStudy (otherworldly):

PEMDAS do u know what that is

OpenStudy (klepez):

\[2(x+2)^2 +3\]

Awolflover1 (awolflover1):

\[2x^2+8x+8+3\]

OpenStudy (klepez):

ya after you distribute the square where do you go

OpenStudy (otherworldly):

its not 8x its 16x

Awolflover1 (awolflover1):

\[(2x^2)+(8x)+(8+3)\]

OpenStudy (otherworldly):

its not (8x) because 2 x 2 = 4 then 4 x 4

OpenStudy (klepez):

\[2(x^2+2^2)+3\]

OpenStudy (klepez):

right?

OpenStudy (otherworldly):

no it would be ( 2x + 4) ^2 +3 then go from there

OpenStudy (klepez):

then I have 2x^2+8+3 or 2x^2+11

OpenStudy (otherworldly):

no its not because e2x^2 is 2x times 2x so 4x^2 + 16 + 3

OpenStudy (otherworldly):

the power of two means double the same number not times 2

OpenStudy (otherworldly):

so it would be 4 times 4 = 16

OpenStudy (klepez):

4x+19

OpenStudy (otherworldly):

yes then 4x= -19

OpenStudy (otherworldly):

find x

jhonyy9 (jhonyy9):

2(x+2)^2 +3 = 2(x^2 +4x +4) +3 = 2x^2 +8x +8 +3 = 2x^2 +8x +11 @Awolflover1 your result is right sure

jhonyy9 (jhonyy9):

so what you need to do - solve it for x ? in this case discriminant = 64-88=-24 <0 what mean that have not real roots hope helped